### v1

# LiV2O4: Hund-Assisted Orbital-Selective Mottness

Abstract
We show that the remarkably small Fermi-liquid coherence scale and large effective mass observed in LiV2O4 are due to the proximity of a Hund-assisted orbital-selective Mott state. Our work is based on an ab initio dynamical mean-field approach, combining several quantum impurity solvers to capture the physics from high to very low temperature. We find that the Hund coupling plays a crucial role in rearranging the orbital populations and in generating the heavy mass and low coherence scale. The latter is found to be approximately 1–2 Kelvin, even though the most correlated orbital is found to be significantly doped ($\sim 10\%$) away from half-filling. A flat quasiparticle band appears near the Fermi level as a result of the strong electronic correlations. Finally, we discuss our results in comparison to experiments.
The ‘heavy fermion’ (HF) phenomenon – quasiparticles acquiring very large effective masses – is one of the most spectacular manifestations of strong electronic correlations. It is usually observed in materials having conduction electrons hybridizing with very localized degrees of freedom, typically $f$-orbitals in rare-earth or actinide compounds. In this context, the discovery of a remarkably large specific-heat enhancement in LiV2O4 (LVO), a transition-metal oxide involving no $f$-electrons, came as quite a surprise [1, 2, 3]. Indeed, only few other transition-metal compounds displaying very large mass enhancements are known to this day [4, 5].
Figure 1: (a) DFT band structure of LVO for the four $a_{1g}$- and eight $e_{g}^{\pi}$-like bands around the Fermi level. The eigenvalues of the MLWFs perfectly match the DFT bands [6]. Color: orbital character. Inset: Brillouin zone with high-symmetry points. (b) DFT DOS and projections onto the MLWFs.
The electronic structure of this material obtained with the density functional theory (DFT) displays two sets of bands close to the Fermi level as shown in Fig. 1 [7, 8]: one set of $a_{1g}$ character, with a narrow bandwidth ($\sim\!0.7$ eV), and one set of $e_{g}^{\pi}$ character with a broader bandwidth ($\sim\!1.9$ eV). A connection to HF materials was therefore suggested in early works [9, 8, 10, 11], with the $a_{1g}$ (resp. $e_{g}^{\pi}$) band playing the role of localized states (resp. conduction electrons). It was later realized, though, that this picture is not tenable. From a theoretical viewpoint, the narrow band is too broad to play the role of localized states, and, importantly, the coupling between the narrow and wide band is dominantly an intra-atomic Hund coupling which favors spin alignment, not an antiferromagnetic Kondo exchange [12, 13, 14]. Furthermore, some experimental features are markedly different from conventional HFs. The onset of Kondo coherence in HFs is typically signaled by a coherence peak in the temperature ($T$) dependence of the resistivity, followed by a lower scale $T_{\mathrm{FL}}\,$below which $\sim T^{2}$ Fermi-liquid (FL) behavior holds. While a small FL scale $T_{\mathrm{FL}}\,$$\sim\!2$ K is indeed observed in LVO, the resistivity, however, monotonously increases upon heating, eventually reaching a ‘bad metal’ regime at high $T$ [2, 3, 15]. From various experimental probes, the onset temperature of electronic coherence is estimated as $T_{\mathrm{onset}}\,$$\sim$ 10–30 K [2, 16, 3].
Aware of the difficulties of the Kondo scenario, Arita et al. proposed an alternative explanation [12]. In their picture, all the action takes place in the narrow $a_{1g}$ band, which is viewed as a doped Mott insulator, while the broader $e_{g}^{\pi}$ bands are merely spectators. In this description, a tiny hole doping ($\sim\!2\%$) of the $a_{1g}$ band yields the observed large effective mass. In support of this picture, the authors of Ref. [12] performed dynamical mean-field theory (DMFT) [17] calculations on a simplified two-orbital model down to room temperature (as well as projective Monte Carlo calculations at $T\!=\!0$). However, because of the very low energy scales involved, a full calculation of a realistic model for LVO in the DFT+DMFT framework covering the whole $T$ range from above $T_{\mathrm{onset}}\,$to below $T_{\mathrm{FL}}\,$has not been achieved yet [18].
In this Letter, we leverage on recent advances in DMFT impurity solvers to achieve this goal and elucidate the physics responsible for the HF behaviour in LVO. To probe the full crossover from a higher-$T$ incoherent metal to the low-$T$ heavy-mass FL and to access the low FL scale, we use a combination of three DMFT solvers: Quantum Monte Carlo (QMC) for 10 K $\!<\!T\!<\!$ 300 K, tensor networks (TN) for $T\!>\!$ 1 K and the numerical renormalization group (NRG) for $T\!<\!$ 1 K.
We show that LVO is close to a Hund-assisted orbital-selective Mott state [19, 20, 21, 14, 5]. The inter-orbital Hund coupling $J$ (and therefore the $e_{g}^{\pi}$ bands) plays a central role. First, it induces a redistribution of the ($a_{1g}$, $e_{g}^{\pi}$) orbital populations per vanadium due to interaction from $(0.4,2\times 0.55)$ in DFT to $(0.9,2\times 0.3)$ in DMFT. Second, it is responsible for the heavy mass and low $T_{\mathrm{FL}}\,$ via the spin-blocking mechanism characteristic of Hund systems in which electrons can only hop between atomic configurations with maximum spin [22, 23, 24, 25]. LVO is shown to be close to an orbital-selective Mott state in which the $a_{1g}$ orbital is prevented to fully localize at $T=0$ due to inter-orbital hopping [26]. Remarkably, we obtain a low $T_{\mathrm{FL}}\,$ already at $\sim\!10\%$ doping of the $a_{1g}$ band, in contrast to the $\sim\!2\%$ doping of Ref. [12].
Electronic structure and effective model—LVO crystallizes in the fcc spinel structure (space group Fd3m). Crystal structure data from neutron scattering are available down to 12 K [1]. The primitive unit cell contains 14 atoms, with 4 V atoms, each embedded in an octahedral crystal field of the surrounding oxygen atoms. The octahedra themselves are corner-sharing and the local point group is trigonal $D_{3d}$, lowering the symmetry of the system and splitting the $t_{2g}$ states into a single $a_{1g}$ and a doubly degenerate $e_{g}^{\pi}$ orbital higher in energy.
We perform DFT calculations using the Quantum ESPRESSO software package fixing the structural parameters to the 12 K experimental values [1]. Maximally localized Wannier functions (MLWFs) for the $a_{1g}$ and $e_{g}^{\pi}$ states are constructed with Wannier90 [27], accurately reproducing the DFT low-energy electronic structure. In Fig. 1(a), the bandstructure of the Wannier Hamiltonian is shown together with its projection on the $a_{1g}$ and $e_{g}^{\pi}$ states. The projected density of states (DOS) in Fig. 1(b) reveals that the $a_{1g}$ DOS is sharply peaked at 0.1 eV above the Fermi level, whereas the $e_{g}^{\pi}$ DOS is relatively flat.
The constrained random phase approximation [28] as implemented in the RESPACK code [29] is used to determine the effective Coulomb interaction at low energy. We focus on the static, $\omega\!=\!0$ limit and fit the resulting four-index tensor to the symmetrized Kanamori form (including spin-flip and pair-hopping terms) with three independent parameters. The optimal fit yields $U\!=\!3.94$ eV, $U^{\prime}\!=\!2.83$ eV, and $J\!=\!0.56$ eV, which violates the full rotational invariance ($U^{\prime}=U\!-\!2J$) by only $\sim 0.01$ eV.
Figure 2: (a) $\mathrm{Im}\Sigma(i\omega_{n})$ for the $a_{1g}$ orbital (circles) and $e_{g}^{\pi}$ orbitals (crosses) at various $T$ from DFT+DMFT using QMC, TN, and NRG. Inset: zoom at low frequencies. (b) Scattering rate and (c) spectral function at the Fermi level as a function of $T$. Results at $i0^{+}$ in (b)–(c) for QMC and TN were obtained by fitting a fourth-order polynomial to the lowest eight Matsubara points. Insets in (b)–(c) magnify the FL behavior.
We use DMFT to solve the effective many-body model, with a combination of impurity solvers to cover all $T$ regimes (see [6] for details). First, we use the continuous-time hybridization-expansion QMC solver implemented in the TRIQS software library [30, 31] and its interface to electronic structure codes TRIQS/DFTTools and TRIQS/solid__\__dmft [32, 33, 6]. QMC calculations were performed down to 11.6 K ($1/1000$ eV). Second, the $T\!=\!0$ TN-based solver from [34, 35, 36] was extended to $T\!>\!0$ using thermal-state purification [37, 38] and applied down to 2.9 K ($1/4000$ eV). Third, to access the sub-$1\,\mathrm{K}$ regime, we use NRG, similarly as in Refs. [39, 40]. To make non-degenerate three-orbital NRG calculations tractable, we increase the local symmetry by neglecting the pair-hopping part of the interaction when using this solver [39, 6].
Crossover to a heavy FL—The emergence of the low-$T$ heavy FL is illustrated in Fig. 2(a). It shows the imaginary part of the self-energy for each orbital as a function of Matsubara frequency for several $T$. The $a_{1g}$ self-energy has a strong $T$ dependence at low frequency, with a sudden decrease below the onset temperature $T_{\mathrm{onset}}\,$$\sim 12$ K. This is a hallmark of strong correlations and corresponds to a sharp crossover from a high-$T$ incoherent regime with a large scattering rate to a low-$T$ coherent regime. This is further supported by Fig. 2(b), which shows the extrapolated zero-frequency value of the $a_{1g}$ scattering rate $-\mathrm{Im}\Sigma_{a_{1g}}(i0^{+})$ as a function of $T$. The local spectral function (DOS) at the Fermi level $A_{a_{1g}}(0)$ correspondingly increases, as shown in Fig. 2(c). By contrast, the $e_{g}^{\pi}$ self-energy is smaller and depends more weakly on temperature and frequency.
The FL fully develops only below the FL coherence temperature $T_{\mathrm{FL}}\,$$\sim$ 1–2 K, an order of magnitude smaller than $T_{\mathrm{onset}}\,$. The FL regime is characterized by $-\mathrm{Im}\Sigma\!\sim\!\omega^{2}\!+\!\pi^{2}T^{2}$ (see End Matter for $T\!=\!0$ and Fig. 2(b) for $\omega\!=\!0$) for the retarded self-energy and $\chi^{\prime\prime}\!\sim\!\omega$ for the imaginary part of the retarded local spin susceptibility (see End Matter). The quasiparticle weights (from NRG at $T\!=\!0$) are $Z_{a_{1g}}\!\approx\!0.003$ and $Z_{e_{g}^{\pi}}\!\approx\!0.03$.
Figure 3: (a) DFT+DMFT (QMC+Padé) $k$-resolved spectral function at $T\!=\!11.6$ K (black lines give the DFT bands). Inset: zoom to the Fermi level, showing a flat quasiparticle band with suppressed spectral weight directly below $\omega\!=\!0$. (b) DOS in the same energy range (see [6] for a wider range). The plain (dashed) blue line refers to the $a_{1g}$ ($e_{g}^{\pi}$) orbital. The $a_{1g}$ result at larger $T$ is also shown. (c) Real and (d) imaginary part of the QMC+Padé real-frequency self-energies.
Quasiparticle bandstructure—Figure 3(a) shows the momentum-resolved spectral function summed over orbitals at $T=11.6$ K. The momentum-integrated spectral functions are displayed in panel (b) for several $T$. These results were obtained by analytic continuation of the QMC self-energies using Padé extrapolation, see Fig. 3(c)–(d), and are consistent with NRG [6].
We see from Fig. 3(b) that the onset of coherence is associated with the growth of a remarkably narrow quasiparticle peak at the Fermi level, in line with the small $Z_{a_{1g}}$. This corresponds to the formation of a flat band near the Fermi level (see panel (a)), together with the strong renormalization of the overall band structure (compare the black lines from DFT). The formation of the flat band can be understood from the fact that $\mathrm{Re}\Sigma_{a_{1g}}(\omega)$ (panel (c)) exhibits a very steep rise near $\omega\!=\!0$. As a result, the quasiparticle equation (quoting for simplicity a one-band version) $\omega-\mathrm{Re}\Sigma(\omega)=\varepsilon_{k}-\mu$ has solutions $\omega_{k}$ which remain very close to zero energy for an extended range of momenta, as seen in panel (a) near the L-point.
Discussion—To elucidate the mechanism responsible for the strong correlations observed above, we first consider the effect of varying the Hund coupling $J$. Figure 4(a) shows the $a_{1g}$ self-energy on the Matsubara axis at $T\!=\!11.6$ K for various $J$. At small $J$, the self-energies are small; for $J\!\gtrsim\!0.3$ eV, strong correlations develop. The self-energy becomes large and acquires the characteristic frequency dependence emphasized above (Fig. 2(a)). This frequency dependence is reminiscent of Hund metals in the ‘spin freezing’ regime [22, 23, 41] (see, e.g., Fig. 8 of Ref. [41]). Correspondingly, the extrapolated zero-frequency scattering rate at finite $T$ quickly increases as a function of $J$ (inset of panel Fig. 4(a)), indicating a suppressed coherence scale. The Hund coupling is known to induce strong correlations through the ‘spin blocking’ mechanism, which forces electrons to keep a high-spin configuration while hopping between different sites. The dominance of high-spin configurations is indeed supported by the valence- and spin-resolved histograms, obtained with QMC, see Fig. 4(b), or NRG [6].
Figure 4: (a) Evolution of $\Sigma_{a_{1g}}$ with increasing $J$ at $T\!=\!11.6$ K. Results are obtained with the TN solver at fixed $U\!=\!3.94$ eV and $U^{\prime}\!=\!U\!-\!2J$. Inset: scattering rate obtained as in Fig. 2(b). (b) Histogram displaying the statistical weights of the relevant multiplets in each valence ($N$) and spin ($S$) sector, obtained from QMC at T=11.6 K. (c) Evolution of the effective crystal field coming from $\mathrm{Re}\Sigma(0)$ and of the orbital populations with increasing $J$. (d) $T_{\mathrm{FL}}\,$for LVO as a function of the $a_{1g}$ occupancy when changing the crystal field (as indicated in the top horizontal scale, see text) and for an effective model involving only the $a_{1g}$ orbital, using the same intra-orbital $U$ value. $T_{\mathrm{FL}}\,$is estimated via the peak position of $\chi^{\prime\prime}(\omega)$ at $T\!=\!0$ from NRG (see End Matter). The vertical lines in panels (c) and (d) indicate the LVO values.
In Fig. 4(c), we show the $a_{1g}$ and $e_{g}^{\pi}$ orbital occupancies as a function of $J$. The Hund coupling induces a significant redistribution of orbital populations [18], with the occupancy of the $a_{1g}$ orbital evolving from the DFT value $\sim\!0.4$ to $\sim\!0.9$ as $J$ is increased beyond $J\!\sim\!0.3$ eV. Hence, the $a_{1g}$ orbital becomes $\sim\!10\%$ hole-doped away from half-filling. This points at the relevance of Mott physics in this material, as proposed in Ref. [12].
To analyze the possible proximity of a Mott-insulating state, we perform a numerical experiment in which we add a crystal field to the Wannier Hamiltonian, while keeping the total $d$-electron count equal to 1.5 by adjusting the chemical potential. Figure 4(d) shows $T_{\mathrm{FL}}\,$determined by NRG as a function of the $a_{1g}$ orbital occupancy (lower horizontal scale) or the applied crystal field (upper scale). We see that $T_{\mathrm{FL}}\,$is driven to remarkably low values as half-filling of the $a_{1g}$ orbital is approached, with values as low as $10^{-6}$ K at $1\%$ doping. This can be interpreted as the system reaching an orbital-selective Mott (OSM) phase [19, 20, 21, 14, 5] in which the $e_{g}^{\pi}$ electrons remain itinerant while the $a_{1g}$ ones basically localize, full localization being prevented at $T\!=\!0$ by inter-orbital hopping [26]. The relevance of the OSM phenomenon and of Hund physics as a possible cause for $d$-electron HF behavior was recently emphasized by Crispino et al. [5] and also previously mentioned qualitatively for LVO in the NMR study of Shimizu et al. [14].
While our results support the relevance of Mottness for LVO [12], they provide novel insights into the mechanism responsible for the HF behavior in LVO, namely the decisive role of Hund coupling and the wide range of permissible fillings. In Fig. 4(d), we compare $T_{\mathrm{FL}}\,$of LVO, as a function of orbital occupancy, to that of a single-band model with the same DOS as the one of the $a_{1g}$ orbital (also solved in DMFT). We see that, in order to reach a value of $T_{\mathrm{FL}}\,$in the range 1–2 K, as observed in experiments, a tiny doping of $\sim\!1\%$ is required in the single-band picture. By contrast, the small $T_{\mathrm{FL}}\,$is robustly generated from the realistic multi-orbital description, as a result of orbital and Hund physics, in a wide range of $a_{1g}$ doping. More broadly, our results emphasize that the multi-orbital character of the system is crucial, with inter-orbital correlations playing a key role, and hence that the $e_{g}^{\pi}$ bands are not merely spectators.
[…]
The low-$T$ specific heat coefficient $C=\gamma T+\dots$ can be evaluated within DMFT from the zero-frequency spectral functions and quasiparticle weights of each orbital $m$ as $\gamma\!=\!\frac{2\pi^{2}}{3}k_{B}^{2}\sum_{m}A_{m}(0)/Z_{m}$. The bare DFT value from our calculated DOS is $\gamma_{\mathrm{DFT}}\!=\!17.1$ mJ/mol K-2, in good agreement with previous theoretical estimates [18, 7]. Reported experimental values in the low-$T$ HF regime are in the range $\gamma/\gamma_{\mathrm{DFT}}\!\sim\!25$ [1, 3] to $\!\sim\!30$ [44]. A comparable value is reached in our DMFT calculations at $T\!\sim\!8$ K, but, at low $T$, we overestimate $\gamma/\gamma_{\mathrm{DFT}}$ by a large factor ($\sim\!10$). We note, however, that no clear saturation of $C/T$ at low $T$ is seen experimentally, and that further increase of $\gamma$ at low $T$ was reported on some samples [44, 45], so that the precise behavior of the specific heat of LVO at low $T$ is yet to be fully clarified.
We conjecture that the overestimation of $\gamma$ at low $T$ comes from the intrinsic limitations of single-site DMFT. In this approach, $\gamma$ is inversely proportional to $Z$, which we find to be tiny for the $a_{1g}$ orbital. In a more accurate treatment beyond the single-site approximation, the inter-site magnetic exchange should intervene and reduce $C/T$ by reducing the entropy associated with fluctuating local moments above $T_{\mathrm{onset}}\,$. Indeed, slave-particle [46, 47] or cluster-DMFT calculations [48] yield, for a single band model, $\gamma/\gamma_{\mathrm{DFT}}\sim 1/(Z+J_{\mathrm{AF}}/\epsilon_{\mathrm{F}})$ with $J_{\mathrm{AF}}$ the antiferromagnetic super-exchange and $\epsilon_{\mathrm{F}}$ a typical (bare) electronic scale. In many oxides, the exchange term limits the value of $\gamma$ as compared to the DMFT $1/Z$ enhancement. Remarkably, this effect appears to be much less pronounced in LVO, allowing $\gamma$ to reach the large observed value. This is most likely due to frustration [44, 49, 50, 51, 14, 52, 45] which results in a large amount of entropy stored at low $T$ and hence a reduced value of the exchange term $J_{\mathrm{AF}}/\epsilon_{\mathrm{F}}$. NMR [53] and neutron scattering [51, 52] experiments yield estimates of the effective exchange between V-atoms in the range $\sim$ 20–30 K, comparable to $T_{\mathrm{onset}}\,$and significantly smaller than typical values in unfrustrated oxides. A proper description of the spin correlations probed by these experiments likely requires a treatment beyond single-site DMFT [6].
Outlook—The physical picture emerging from our work is that LVO is close to a Hund-assisted orbital selective Mott state. Recent experiments have indeed revealed that an insulating state can be induced by Li-intercalation [54] or, together with charge ordering, when subjecting this compound to uniaxial strain [55]. While some qualitative aspects of this physical picture were anticipated before [12, 14, 5], our work demonstrates how key progress on DMFT solvers now allows us to cover the full range of energy scales in this particularly challenging case and reliably establish the physical mechanism at hand. Our work also points at two necessary further developments: (i) including spatial fluctuations and computing spin and charge responses beyond single-site DMFT and (ii) exploring instabilities toward charge and spin ordering dependent on external perturbations such as pressure, uniaxial strain, and chemical substitutions.
Acknowledgements.
Acknowledgements—We acknowledge useful discussions and communications with Ryotaro Arita, Sophie Beck, Riccardo Comin, Andrea Damascelli, Luca de’ Medici, Dongjin Oh, Dennis Huang, Xiangyu Luo, Giorgio Sangiovanni, Hidenori Takagi, Manish Verma, and Jan von Delft. M.G. and U.S. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-426 EXC-2111-390814868. M.G. and U.S. are grateful to the Flatiron Institute and A.G. to the University of Munich for their hospitality. The Flatiron Institute is a division of the Simons Foundation.